# Flyby-induced displacement: analytic solution

### v1

@@ Abstract
Approximating the derivative-of-the-Gaussian profile proposed by Gibbons and Hawking by the scarf potential, the scattering of particles by a gravitational wave generated by flyby is described analytically by following the Nikiforov-Uvarov method. Pure displacement arises when the wave zone contains an integer number of half-waves. The results confirm the prediction of Zel’dovich and Polnarev.
@@ I Analytic solution for flyby
Zel’dovich and Polnarev suggested that gravitational waves generated by flyby could be observed through the displacement memory effect (DM) ZelPol ; DM-1 ; Jibril . Since then, and with the discovery that gravitational memory effects are closely related to soft graviton theorems and Bondi-Metzner-Sachs symmetries at future null infinity, the field has grown considerably BMS1 ; BMS2 ; BMS3 ; BMS4 . Zel’dovich - Polnarev suggestion was confirmed numerically for the derivative-of-the-Gaussian profile ${\cal A}^{dG}\propto d\big{(}\exp[-U^{2}]\big{)}/dU$ proposed by Gibbons and Hawking GibbHaw71 provided the parameters take some “magic” values DM-1 . Their profile does not support an analytic solution, however it can be approximated by the derivative of the Pöschl-Teller potential PTeller ; Chakra ; DM-1 ,
@@
which can be studied semi-analytically using confluent Heun functions DM-1 ; DM-2 . This led us to conjecture a sort of quantisation condition: DM ariseses when the wave zone contains an integer number $m$ of half-waves. In this note we present yet another approximation of the flyby profile, namely by the Scarf potential Scarfpot ,
@@
Note that ${\cal A}^{Scarf}(U)\propto\cosh U\,{\cal A}^{dPT}(U)$ implying that for small values of $U$ it behaves as (1) but for large $U$ it falls off more slowly. However dilating $U$ appropriately, almost perfect overlapping is obtained, as depicted in FIG.1.
Figure 1: The (dilated) Scarf potential (2) and its Pöschl-Teller  alter-ego (1) are both good approximations of the derivative-of-Gaussian profile proposed by Gibbons and Hawking GibbHaw71 to describe flyby.
@@
Below we show that the hyperbolic Scarf potential Scarfpot ; Scrf1 , eqn. (2), allows us to find an analytic solution which confirms the quantisation conjecture above. Eqn.(3) can be viewed indeed as a Schrödinger equation for the (non-normalizable) zero-energy ground state DM-1 . The solution is obtained by switching to $t=\sinh{U}$ which takes (3) into  [^4] For comparison, putting $t=\tanh U$ carries (3) with Pöschl-Teller profile $(g/2)\cosh^{-2}U$ into $(1-t^{2})\frac{d^{2}X}{dt^{2}}-2t\frac{dX}{dt}-\frac{g}{4}X=0\,,$ which can be solved by Legendre polynomials DM-1 .
@@
The polynomial $y_{n}(t)$ is given by the Rodrigues formula Eq.(36) in which we need to find the weight function $\rho(t)$, for which we have the equation,
$\frac{((1+t^{2})\rho)^{\prime}}{(1+t^{2})\rho}=\frac{\tau}{\sigma}=-\frac{(2n-1)t+2\,\mathrm{sign}(g)\sqrt{n(n+1)}}{1+t^{2}}\,,$ (21)
which can be again easily integrated with the result,
@@
The normalization constants $B_{n}$ should be determined from the initial conditions. The first three polynomials are $y_{1}(t)\sim\tau(t)=-t-2\sqrt{2}\,\mathrm{sign}(g)$, $y_{2}(t)\sim 2t^{2}+8\sqrt{6}\,\mathrm{sign}(g)t+23$ and $y_{3}(t)\sim-6t^{3}-72\sqrt{3}\,\mathrm{sign}(g)t^{2}-423t-172\sqrt{3}\,\mathrm{sign}(g)$. These polynomials can be expressed in terms of Routh-Romanovski polynomials Scarfpot ; Raposo as follows: $y_{n}(t)=R_{n}^{(\alpha,\beta)}(t)$, with $\alpha=-2\,\mathrm{sign}(g)\sqrt{n(n+1)}$ and $\beta=-n+\frac{1}{2}$.
@@
Further physical applications of the Scarf potential which include supersymmetry Scarfpot ; Compean , non-central potentials Dutt and even quarks (QCD), are reviewed in ref. Raposo . Note that, in the supersymmetry literature, the Scarf hyperbolic potential, which has recently attracted much attention Scrf1 , was originally called the Gendenstein potential Natanson .

### v2

@@ Abstract
Approximating the derivative-of-the-Gaussian profile proposed by Gibbons and Hawking by the hyperbolic Scarf potential, the scattering of particles by a gravitational wave generated by flyby is described analytically by following the Nikiforov-Uvarov method applied to the Scarf potential. Pure displacement arises when the wave zone contains an integer number of half-waves. The results confirm the prediction of Zel’dovich and Polnarev.
@@ I Analytic solution for flyby
Zel’dovich and Polnarev suggested that gravitational waves generated by flyby could be observed through the displacement memory effect (DM) ZelPol ; DM-1 ; Jibril . Since then, and with the discovery that gravitational memory effects are closely related to soft graviton theorems and Bondi-Metzner-Sachs symmetries at future null infinity, the field has grown considerably BMS1 ; BMS2 ; BMS3 ; BMS4 . Zel’dovich - Polnarev suggestion was confirmed numerically for the derived-of-the-Gaussian profile ${\cal A}^{dG}\propto d\big{(}\exp[-U^{2}]\big{)}/dU$ proposed by Gibbons and Hawking GibbHaw71 provided the parameters take some “magic” values DM-1 . Their profile does not support an analytic solution, however it can be approximated by the derived Pöschl-Teller potential PTeller ; Chakra ; DM-1 ,
@@
which can be studied semi-analytically using confluent Heun functions DM-1 ; DM-2 . This led us to conjecture a sort of quantisation condition: DM ariseses when the wave zone contains an integer number $m$ of half-waves. In this note we present yet another approximation of the flyby profile, namely by the hyperbolic Scarf potential scarfpot ,
@@
For small values of $U$ it behaves as (1) but for large $U$ it falls off more slowly. However dilating $U$ appropriately almost perfect overlapping is obtained, as depicted in FIG.1.
Figure 1: The derived Pöschl-Teller  profile (1) and its Scarf potential alter-ego (2) are both good approximations of the derived-Gaussian one proposed by Gibbons and Hawking GibbHaw71 to describe flyby.
@@
Below we show that the hyperbolic Scarf potential scarfpot ; Scrf1 , eqn. (2), allows us to find an analytic solution which confirms the quantisation conjecture above. Eqn.(3) can be viewed indeed as a Schrödinger equation for the (non-normalizable) zero-energy ground state DM-1 . The solution is obtained by switching to $t=\sinh{U}$ which takes (3) into  [^4] For comparison, putting $t=\tanh U$ carries (3) with Pöschl-Teller profile ${\cal A}^{PT}=(g/2)\cosh^{-2}U$ into a form which can be solved by Legendre polynomials DM-1 .
@@
The polynomial $y_{n}(t)$ is given by the Rodrigues formula Eq. (35) in which we need to find the weight function $\rho(t)$, for which we found,
@@
The normalization constants $B_{n}$ should be determined from the initial conditions. The first three polynomials are $y_{1}(t)\sim\tau(t)=-t-2\sqrt{2}\,\mathrm{sign}(g),\;y_{2}(t)\sim 2t^{2}+8\sqrt{6}\,\mathrm{sign}(g)t+23,\;y_{3}(t)\sim-6t^{3}-72\sqrt{3}\,\mathrm{sign}(g)t^{2}-423t-172\sqrt{3}\,\mathrm{sign}(g)\,.$ They can be expressed in terms of Routh-Romanovski polynomials scarfpot ; Raposo as $y_{n}(t)=R_{n}^{(\alpha,\beta)}(t)$ with $\alpha=-2\,\mathrm{sign}(g)\sqrt{n(n+1)}$ and $\beta=-n+\frac{1}{2}$.
@@
Further physical applications of the Scarf potential which include supersymmetry scarfpot ; Compean , non-central potentials Dutt and even quarks (QCD), are reviewed in ref. Raposo . Note that, in the supersymmetry literature, the Scarf hyperbolic potential has recently attracted much attention Scrf1 . This potential was discovered by Gendenshtein and is therefore referred also to as the Gendenstein potential Gendenon ; Natanson . See also Hait ; Bhattacharya23 ; Bhattacharya:2025ljc for related work.
For isotropic profile we have a “universal model” : every plane gravitational wave can be mapped locally conformally to free Minkowski space by following Niederer, or Arnold Niederer ; Arnold ; Takagi , consistently with the vanishing of the Cotton or Weyl tensor Inomata ; Dhasmana ; GibbonsDark ; ZZH21 .
However gravitational waves are isotropic only in 1d, and their anisotropy in d=2 dim implies that we arrive to unchartered territory Lukash ; Inomata ; ZZH21 and it is not clear whether we still have a (conformal) “universal ancestor”.